# Rice-like complexity lower bounds for Boolean and uniform automata networks

### v1

@@ 2.4 Tree-decompositions and treewidth
    For each vertex $x$ of $G$, the subgraph $T^{\prime}$ of $T$ restricted to vertices whose bag contains $x$ is connected.
A graph may admit multiple tree-decompositions. The width of a tree-decomposition $T$ is the size of its largest bag minus one, ie. $w(T)=\max\{|B(t)|\mid t\in V(T)\}-1$. The treewidth of a graph $G$, written $tw(G)$, is the minimal width among all of its tree-decompositions. A tree-decomposition $T$ of width $w(T)=k$ is called a $k$-tree-decomposition, and we assume without loss that all bags of $T$ are of size $k+1$. For the corresponding graph $G$, it means that $tw(G)\leq k$. Remark that a functional digraph can be very similar to its tree-decomposition.
Given a graph $G$, a tree-decomposition $T$ and an MSO formula $\psi$, for each $v\in V(T)$ we denote:
  • $\mathcal{S}_{T}(v)$ the largest subtree of $T$ rooted in $v$,
  • $\mathcal{N}_{T}(v)$ the boundaried graph resulting from the subgraph of $G$ spanned by the vertices appearing in bags of $\mathcal{S}_{T}(v)$, ie. $\bigcup_{u\in V(\mathcal{S}_{T}(v))}B(u)$, with ports $B(v)$,
  • $\mathcal{C}_{T}(v)$ is the equivalence class of $\mathcal{N}_{T}(v)$ for the relation $\sim_{\psi}$.
See Figure 4 for an example.
Figure 4: A graph $H$ (left) and its tree-decomposition $T$ (middle). In the tree-decomposition a node $v$ such that $B(v)=(4,5)$ is highlighted, with the subtree $\mathcal{S}_{T}(v)$ surrounded with a dashed line. The graph $\mathcal{N}_{T}(v)$ (right) where the primary ports, from the root bag of $\mathcal{S}_{T}(v)$, are hatched.
When the tree $T$ is clear from the context, we drop the subscript notation. Two nodes $v,v^{\prime}\in V(T)$ are said to be $\mathcal{N}$-different, written $v\neq v^{\prime}$, when $\mathcal{N}(v)\neq\mathcal{N}(v^{\prime})$. The $\mathcal{N}$-size of $T$ is its number of $\mathcal{N}$-different nodes (hence at most $2^{|G|}$, because the order of vertices inside the bags does not matter). The $\mathcal{N}$-length of a path $\mathcal{P}$ in $T$ is the number of $\mathcal{N}$-different nodes it goes through.
When gluing two $k$-tree-decompositions $T\oplus T^{\prime}$, ports will be single vertices (in the trees, $P_{1}(T^{\prime})$ will be the root of $T^{\prime}$ and $P_{2}(T)$ will be a leaf of $T$) and we merge their bags (of size $k+1$) by renaming all the bags of $T^{\prime}$ according to the substitution $B(P_{1}(T^{\prime}))\leftarrow B(P_{2}(T))$. Formally, we define $T\oplus T^{\prime}=T^{\prime\prime}$ in terms of $1$-graphs, and set the following bags for $T^{\prime\prime}$:
  • For all $v\in V(T)$, set $B_{T^{\prime\prime}}(v)=B_{T}(v)$,
  • For all $v\in V(T^{\prime})\setminus P_{1}(T^{\prime})$, set $B_{T^{\prime\prime}}(v)=B_{T^{\prime}}(v)[B(P_{1}(T^{\prime}))\leftarrow B(P_{2}(T))]$.
In words, we merge the bags of the merged vertices, and propagate this substitution to all the other bags of $T^{\prime}$. See Figure 5 for an example.
Figure 5: Graphs $G$, $G^{\prime}$ and $G\oplus G^{\prime}$ from Figure 3, with their respective tree-decompositions below. In the trees, primary ports are represented by hatched bags, and secondary ports by grey bags. Let $T$ (left) and $T^{\prime}$ (middle) be the respective tree-decompositions of $G$ and $G^{\prime}$, then $T\oplus T^{\prime}$ (right) is a tree-decomposition of $G\oplus G^{\prime}$. Notice the propagation of the renaming of vertices from the merged bags: $(7,6)\leftarrow(4,5)$.
As $\Delta^{\Gamma}(\omega)$ for graphs, we define a similar operation for $k$-tree-decompositions. Let $\mathcal{T}=\{T_{i}\}_{i\in\mathcal{I}}$ be a finite family of $k$-tree-decompositions, each with their root as primary port, and a leaf as secondary port. For $\omega=\omega_{1}...\omega_{\mu}$ a nonempty word over alphabet $\mathcal{I}$, the operation $\Lambda^{\mathcal{T}}$ is defined by induction on $\mu$ as:
$\Lambda^{\mathcal{T}}(\omega_{1})=T_{\omega_{1}},\qquad\Lambda^{\mathcal{T}}(\omega_{1}...\omega_{\mu})=\Lambda^{\mathcal{T}}(\omega_{1}...\omega_{\mu-1})\oplus T_{\omega_{\mu}}.$
Given a formula $\psi$, a $k$-tree-decomposition of $G$ can be viewed as a $\Sigma_{k+1}$-labeled tree, by labeling each node $v$ with $\mathcal{C}(v)$ instead of the bag $B(v)$. This is especially useful thanks to the following remark.
Remark 2 ([10, Remark 15]).
Let $\Gamma=\{G_{i}\}_{i\in\mathcal{I}}$ be a family of $k$-graphs and $\mathcal{T}=\{T_{i}\}_{i\in\mathcal{I}}$ be a family of $\Sigma_{k}$-labeled trees, both indexed by the same finite set $\mathcal{I}$. Suppose that for every $i$, the tree $T_{i}$ corresponds to a decomposition of $G_{i}$ such that:
  • the primary port of $T_{i}$ is its root $r$, and $B(r)=P_{1}(G_{i})$,
  • the secondary port of $T_{i}$ is a leaf $\ell$, and $B(\ell)=P_{2}(G_{i})$.
Then, for every nonempty word $\omega$ over alphabet $I$, the tree $\Lambda^{\mathcal{T}}(\omega)$ corresponds to a decomposition of $\Delta^{\Gamma}(\omega)$.
For a formula $\psi$, we say that a $\Sigma$-labeled tree is positive when it corresponds to a model of $\psi$ (and negative otherwise). A $\Sigma$-labeled tree can correspond to several different graphs, but from Remark 2 they are all models, or they are all counter-models.
@@ 3 Previous works
The computational complexity of questions on the dynamics of automata networks has traditionally raised a great interest, as a mean to understand the possibilities and limits of algorithmic problem solving, in this model and multiple variants with restrictions on the type of local functions (e.g. Boolean, theshold, disjunctive), the architecture of the network, or with different update modes. Alon noticed in [1] that it is ${\mathsf{NP}}$-hard to decide whether a given BAN has a fixed point. Floréen and Orponen then settled the complexity of multiple problems related to fixed points [7, 27]. Further developments [17] include the study of limit cycles [3, 31], update modes [28, 26, 3, 2] and specific rules [12, 11, 13].
In a recent series of two papers [9, 10], Rice-like complexity lower bounds have been established, encompassing at once many results from the literature. They state that non-trivial formulas yield algorithmically hard $\psi$-dynamics problems. The first article concentrates on the deterministic setting and FO formulas (using Hanf-Gaifman’s locality), while the second article brings the result to MSO on non-deterministic automata networks (and introduces the notion of arborescent formula).
Theorem 1.a for FO ([9, Theorem 5.2]).
@@
For any arborescent MSO formula $\psi$, problem $\psi$-NAN-dynamics is ${\mathsf{NP}}$-hard or ${\mathsf{coNP}}$-hard.
In the present work, we bring these results to $q$-uniform ANs and NANs, for MSO formulas (Theorem 1). Indeed, the reductions presented in the proofs of [9, Proposition 5.1.2 and Lemma 5.1.4] and [10, Proposition 19 and Lemma 20] may require arbitrary alphabet sizes in the construction of instances of $\psi$-AN-dynamics and $\psi$-NAN-dynamics, respectively. The proof of Theorem 1 is based on the abstract pumping technique introduced in [9] and developed in [10], which is reviewed in the rest of this section.
Let us start with the example of Figure 1. It implements a reduction from SAT (here with $S=[(x_{1}\vee x_{2})\wedge(\neg x_{1}\vee\neg x_{2})\wedge x_{3}]$ on 3 Boolean variables), to prove the ${\mathsf{NP}}$-hardness of the fixed point existence problem ($\exists x:x\to x$). Indeed, for each valuation $v\in\{0,1\}^{3}$ of $S$, the pair of configurations $x,x^{\prime}\in\{0,1\}^{4}$ with $x_{\{1,2,3\}}=x^{\prime}_{\{1,2,3\}}=v$ creates either:
@@
Observe that this same reduction also proves the ${\mathsf{coNP}}$-hardness of asking whether the dynamics is a union of limit cycles of length $2$, and of asking whether the dynamisc is injective, both by reduction from UNSAT. The basic idea to reduce from SAT or UNSAT is to evaluate $S$ on the configuration, and produce (glue) one of two graphs $G_{0},G_{1}$ (using a distinct subset of configurations for each valuation) according to whether the formula is satisfied or not. If the formula is satisfiable then $G_{0}$ appears at least once, otherwise there are only copies of $G_{1}$. The crucial part is then to obtain appropriate graphs $G_{0}$ and $G_{1}$ for any non-trivial (or arborescent) MSO formula $\psi$, and this requires tools from finite model theory (giving additional parts $G_{2},G_{3}$ to deal with). Analogously, the example of Figure 2 is strongly connected if and only if $S$ is a tautology.
Remark that $\psi$ is fixed in the problem definition, therefore $G_{0},G_{1}$ are considered constant. Also, reductions produce circuits for the local functions, not the transition digraph itself. From [10] we first introduce the concept of saturating graph $G_{0}$ (Subsection 3.1), then the requirement on $G_{1}$ to produce a metareduction (Subsection 3.2), and finaly how $G_{1}$ has been obtained via the abstract pumping technique (Subsection 3.3). We close this section with an overview of our contributions on strengthening this proof scheme (Subsection 3.4).
@@ 3.1 Saturating graph
In [10] it is shown that for all $m\in\mathbb{N}$, there exists a saturating graph $\Omega_{m}$ such that, for all MSO formulas $\psi$ of rank $m$, if a graph contains a copy of $\Omega_{m}$ then it is forced to be a model of $\psi$ (if $\Omega_{m}$ is a sufficient subgraph for $\psi$) or a counter-model of $\psi$ (if $\Omega_{m}$ is a forbidden subgraph for $\psi$).
@@
It will be clear from the context which of the two $\Omega_{m}$ we are considering: Proposition 2 for non-deterministic dynamics, and Corallary 3 for deterministic dynamics.
3.2 Metareduction
For a given MSO formula $\psi$ of rank $m$, whether $\Omega_{m}$ is sufficient (a model of $\psi$) or forbidden (a counter-model of $\psi$) indicates whether we respectively reduce from SAT (${\mathsf{NP}}$-hardness) or UNSAT (${\mathsf{coNP}}$-hardness). Given a propositional formula $S$ on $s$ variables, the configuration space $X$ is partitioned into $2^{\ell}$ subsets of configurations, so that $S$ is evaluated on each subset:

### v2

@@ 2.3 Clique-decompositions and cliquewidth
    $\mathrm{recolor}_{f}$ (arity 1): given a recoloring function $f:C\to C$ and a $k$-colored graph $H$, it is the graph $G=(V(H),E(H),C_{G})$ such that $C_{G}=f\circ C_{H}$;
  4. 4.
    $\mathrm{join}_{M}$ (arity 2): given a set $M\subseteq C\times C\times\{R,L\}$ and two $k$-colored graphs $H$ and $H^{\prime}$, it is the graph $G=(V(H)\sqcup V(H^{\prime}),E(H)\sqcup E(H^{\prime})\sqcup E_{R}\sqcup E_{L},C_{G})$, where $E_{R}=\{(u,v)\mid(i,j,R)\in M,u\in H,C_{H}(u)=i,v\in H^{\prime},C_{H^{\prime}}(v)=j\}$ which are the arrows going from vertices of color $i$ in $H$ to vertices of color $j$ in $H^{\prime}$ for triples with $R$ (right, from $H$ to $H^{\prime}$), and symmetrically for $E_{L}=\{(u,v)\mid(i,j,L)\in M,u\in H^{\prime},C_{H^{\prime}}(u)=i,v\in H,C_{H}(v)=j\}$ with $L$ (left, from $H^{\prime}$ to $H$). Moreover, $C_{G}(v)=C_{H}(v)$ if $v\in H$ and $C_{G}(v)=C_{H^{\prime}}(v)$ otherwise.
Let $G$ be a $k$-colored graph. A clique-decomposition $\mathcal{C}_{G}$ of $G$ is a binary tree such that every node is labeled by one of these four operations, and the graph obtained at the root of $\mathcal{C}_{G}$ is $G$ (the leaves of $\mathcal{C}_{G}$ are labeled by constant or constant, they are the vertices of $G$; nodes labeled by recolor have $1$ child; and nodes labeled by join have $2$ children). We say that $\mathcal{C}_{G}$ generates the $k$-colored graph $G$, and that $\mathcal{C}_{G}$ is a clique-decomposition of a (non-colored) graph $G$ if there exists a coloring $C_{G}$ of the graph $G$ such that $\mathcal{C}_{G}$ is a clique-decomposition of $(V(G),E(G),C_{G})$. Observe that a clique-decomposition generates exactly one graph, hence we sometimes identify a clique-decomposition with the graph it generates (in particular, $|\mathcal{C}_{G}|$ will denote its number of constant nodes).
A marked clique-decomposition is a clique-decomposition with a unique marked node where one child is replaced by a node of arity 0 labeled by $\square$. The marked node acts as a join (disjoint union). For a marked clique-decomposition $\mathcal{C}$ and a clique-decomposition $\mathcal{C^{\prime}}$, we denote $\mathcal{C}\oplus\mathcal{C^{\prime}}$ the operation connecting them by replacing the $\square$ child of the marked node of $\mathcal{C}$ with the root of $\mathcal{C^{\prime}}$. We call this operation gluing $\mathcal{C}$ and $\mathcal{C^{\prime}}$. The operation works the same when $\mathcal{C^{\prime}}$ is also marked and in this case $\mathcal{C}\oplus\mathcal{C^{\prime}}$ is marked too, its marked node being the marked node of $\mathcal{C^{\prime}}$. For two clique-decompositions $\mathcal{C}$ and $\mathcal{C^{\prime}}$, we denote $\mathcal{C}\sqcup\mathcal{C^{\prime}}$ the clique-decomposition $\mathrm{join}_{\emptyset}(\mathcal{C},\mathcal{C^{\prime}})$, generating the disjoint union of the two graphs they respectively generate.
The width of a clique-decomposition is the number of colors used during the generation of the colored graph (i.e. not necessarily only the colors present in the graph eventually generated at the root). The cliquewidth of a graph $G$ is the minimal width of a clique-decomposition generating $G$. We say that a graph $G$ has a $k$-clique-decomposition when $G$ can be generated by a clique-decomposition of width at most $k$.
Let $\Gamma=\{\mathcal{C}_{i}\}_{i\in I}$ be a family of clique-decompositions ($\Gamma$ will always be a finite set), where only one $\mathcal{C}_{j}$ for some $j\in I$ is not marked; if $w=w_{1}\dots w_{\ell}$ is a nonempty word over alphabet $I$ where the only occurrence of $j$ is its last letter $w_{\ell}=j$, then define $\Delta^{\Gamma}(w)$ by induction over the length of $w$ as follows:
$\Delta^{\Gamma}(w_{1})=\mathcal{C}_{w_{1}},\hskip 28.45274pt\Delta^{\Gamma}(w_{1}\dots w_{n})=\Delta^{\Gamma}(w_{1}\dots w_{n-1})\oplus\mathcal{C}_{w_{n}}.$
See Figure 4. Observe that $\oplus$ is not commutative. In the case where $\Gamma$ is a set of graphs, by an abuse of notation, we consider $\Delta^{\Gamma}(w)$ to be the graph generated by $\Delta^{\Gamma^{\prime}}(w)$ where $\Gamma^{\prime}$ is the set of clique-decompositions generating $\Gamma$.
Remark 2.
Because of the recoloring functions, when taking two different nodes in a clique-decomposition, the two clique-decompositions consisting of the descendants of the two respective nodes might generate the same graph. Hence, in this article, when we say that two nodes $v$ and $v^{\prime}$ are different in a clique-decomposition or when writing $v\neq v^{\prime}$, we mean that the two clique-decompositions respectively consisting of the descendants of $v$ and the descendants of $v^{\prime}$ do not have the same set of constant nodes (hence the graph have different sets of vertices). In particular, if $v^{\prime}$ is itself a descendant of $v$, the clique-decomposition of the descendants of $v$ contains at least one constant leaf which is not a descendant of $v^{\prime}$.
2.4 Types
Let $G$ be a $k$-colored graph, $a_{1},\dots,a_{p}$ be vertices of $G$ and $A_{1},\dots,A_{q}$ be subsets of vertices of $G$. The quantifier rank $m$ type of $G,a_{1},\dots,a_{p},A_{1},\dots,A_{q}$ is the set of all MSO formulas with $p$ free point variables and $q$ free set variables and quantifier rank at most $m$ which are satisfied by $G,a_{1},\dots,a_{p},A_{1},\dots,A_{q}$. A set of formulas is a realized $p,q$-type of quantifier rank $m$ if it is such a type. The following lemma is an essential key in previous works and the present article:
Lemma 2 ([21, Proposition 7.5]).
Given $p,q$ and $m$, there are finitely many realized $p,q$-types of quantifier rank $m$.
We denote $G,a_{1},\dots,a_{p},A_{1},\dots,A_{q}\equiv_{q}H,b_{1},\dots,b_{p},B_{1},\dots,B_{q}$ when the corresponding types are the same. Our main focus will be on $0,0$-types which we simply call types. Let $T_{m}$ denote the (finite) set of realized types of quantifier rank $m$ (it will always be clear from the context what is the fixed number $k$ of colors). We also attribute a type to each clique-decomposition (the type of the graph it generates), and to each node $v$ of a clique-decomposition (the type of the graph generated by the subtree whose root is $v$).
@@ 3 Previous works
The computational complexity of questions on the dynamics of automata networks has traditionally raised a great interest, as a mean to understand the possibilities and limits of algorithmic problem solving, in this model and multiple variants with restrictions on the kind of local functions (e.g. Boolean, threshold, disjunctive), the architecture of the network, or with different update modes. Alon noticed in [1] that it is $\mathsf{NP}$}-hard to decide whether a given BAN has a fixed point. Floréen and Orponen then settled the complexity of multiple problems related to fixed points [7, 29]. Further developments [19] include the study of limit cycles [3, 33], update modes [30, 28, 3, 2] and specific rules [13, 12, 14].
In a recent series of two papers [11, 10], Rice-like complexity lower bounds have been established, encompassing at once many results from the literature. They state that non-trivial formulas yield algorithmically hard $\psi$-dynamics problems. The first article concentrates on the deterministic setting and FO formulas (using Hanf-Gaifman’s locality), while the second article brings the result to MSO on non-deterministic automata networks (and introduces the notion of cw-non-trivial formula).
Theorem 1.a for FO ([11, Theorem 5.2]).
@@
For any cw-non-trivial MSO formula $\psi$, problem $\psi$-NAN-dynamics is $\mathsf{NP}$}-hard or $\mathsf{coNP}$}-hard.
In the present work, we bring these results to $q$-uniform ANs and NANs, for MSO formulas (Theorem 1). Indeed, the reductions presented in the proofs of [11, Proposition 5.1.2 and Lemma 5.1.4] and [10, Proposition 11 and Proposition 13] may require arbitrary alphabet sizes in the construction of instances of $\psi$-AN-dynamics and $\psi$-NAN-dynamics, respectively. The proof of Theorem 1 is based on the abstract pumping technique introduced in [11] and developed in [10], which is reviewed in the rest of this section, including novel developments to establish a more precise construction for the purpose of further developments in the remaining sections of the present work.
Let us start with the example of Figure 1. It implements a reduction from SAT (here with $S=(x_{1}\vee x_{2})\wedge(\neg x_{1}\vee\neg x_{2})\wedge x_{3}$ on 3 Boolean variables), to prove the $\mathsf{NP}$}-hardness of the fixed point existence problem ($\exists x:x\to x$). Indeed, for each valuation $v\in\{0,1\}^{3}$ of $S$, the pair of configurations $x,x^{\prime}\in\{0,1\}^{4}$ with $x_{\{1,2,3\}}=x^{\prime}_{\{1,2,3\}}=v$ creates either:
@@
Observe that this same reduction also proves the $\mathsf{coNP}$}-hardness of asking whether the dynamics is a union of limit cycles of length $2$, and of asking whether the dynamics is injective, both by reduction from UNSAT. The basic idea to reduce from SAT or UNSAT is to evaluate $S$ on the configuration, and produce (glue with $\oplus$) one of two graphs $G_{0},G_{1}$ (using a distinct subset of configurations for each valuation) according to whether the formula is satisfied or not. If the formula is satisfiable then $G_{0}$ appears at least once, otherwise there are only copies of $G_{1}$. The crucial part is then to obtain appropriate graphs $G_{0}$ and $G_{1}$ for any non-trivial (or cw-non-trivial) MSO formula $\psi$, and this requires tools from finite model theory (giving additional parts $G_{2},G_{3}$ to deal with). Analogously, the example of Figure 2 is strongly connected if and only if $S$ is a tautology.
Remark that $\psi$ is fixed in the problem definition, therefore $G_{0},G_{1}$ are considered constant. Also, reductions produce circuits for the local functions, not the transition digraph itself. From [10] we first introduce the concept of saturating graph $G_{0}$ (Subsection 3.1), then how $G_{1}$ has been obtained via the abstract pumping technique that verify the requirements for the next subsection (Subsection 3.2) and finally the metareduction (Subsection 3.3). We close this section with an overview of our contributions on strengthening this proof scheme (Subsection 3.4).
@@ 3.1 Saturating graph
In [10] it is shown that for all $m\in\mathbb{N}$, there exists a saturating graph $\Omega_{m}$ such that, for all MSO formulas $\psi$ of rank $m$, if a graph contains a copy of $\Omega_{m}$ then it is forced to be a model of $\psi$ (then we say that $\Omega_{m}$ is a sufficient subgraph for $\psi$) or a counter-model of $\psi$ (then we say that $\Omega_{m}$ is a forbidden subgraph for $\psi$).
@@
It will be clear from the context whether we consider $\Omega_{m}$ with out-degree $1$: Proposition 3 for non-deterministic dynamics, and Corollary 4 for deterministic dynamics.
Technically speaking, $\Omega_{m}$ is a fixed graph with some fixed cliquewidth $k^{\prime}$. When taking the disjoint union with $G$ of cliquewidth at most $k$, the resulting graph $G\sqcup\Omega_{m}$ has cliquewidth at most $\max\{k,k^{\prime}\}$. In the sequel, the value of $k$ will be given by the hypothesis of cw-non-triviality on the fixed formula $\psi$ under study, and we will consider without loss of generality that $k\geq k^{\prime}$ so that $G\sqcup\Omega_{m}$ has cliquewidth at most $k$. This detail will be recalled when needed.
3.2 Pumping models of cw-non-trivial sentences
We pump in some clique-decompositions and construct a family of model clique-decompositions. Recall that since there is exactly one graph that corresponds to one clique-decomposition, the construction of model graphs follows. Let us sketch how three graphs are obtained in [10] in order to construct an infinite family of models of $\psi$:
Proposition 5 ([10, Proposition 10]).
Let $\psi$ be an MSO sentence and $k$ an integer. If $\psi$ has infinitely many models of cliquewidth at most $k$, then there exist clique-decompositions $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ such that $\mathcal{C}_{2}\oplus\mathcal{C}_{1}\oplus\dots\oplus\mathcal{C}_{1}\oplus\mathcal{C}_{3}$ is a model of $\psi$, whatever the number of occurrences of $\mathcal{C}_{1}$. Furthermore, $\mathcal{C}_{1}$ contains at least one constant node (i.e., it is not the trivial marked clique-decomposition with a unique node labeled by $\square$).
This proposition relies on the key concept of compositionality, stated as follows (recall that $T_{m}$ denotes the set of realized types of quantifier rank $m$ for $k$-colored graphs):
Lemma 6 ([10, Lemma 9]).
For every marked clique-decomposition $\mathcal{C}$ of width $k$ and every integer $m$, there is a map $\Lambda_{\mathcal{C}}:T_{m}\to T_{m}$, such that for every clique-decomposition $\mathcal{C}^{\prime}$ of type $t$, the type of $\mathcal{C}\oplus\mathcal{C}^{\prime}$ is $\Lambda_{\mathcal{C}}(t)$.
Lemma 6 means that when we have a marked clique-decomposition $\mathcal{C}$ and two clique-decompositions $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ of the same type, then $\Lambda_{\mathcal{C}}(\mathcal{C}_{1})=\Lambda_{\mathcal{C}}(\mathcal{C}_{2})$. In other words, $\mathcal{C}\oplus\mathcal{C}_{1}$ and $\mathcal{C}\oplus\mathcal{C}_{2}$ have the same type. The proof of Proposition 5 is illustrated on Figure 4: any large enough clique-decomposition contains two vertices ($v$ and $v^{\prime}$) of the same type (because $T_{m}$ is finite), consequently a piece of the clique-decomposition ($\mathcal{C}_{1}$) can be copied (pumped) any number of times while preserving the type of the obtained clique-decomposition (by Lemma 6). By definition the type determines whether it is a model or a counter-model of $\psi$, hence Proposition 5 follows.
[…]
Figure 3: Illustration of $\Delta$.
$\mathcal{C}_{2}\bulletvv$$\mathcal{C}_{1}\bulletvv^{\prime}$$\mathcal{C}_{3}$
$\to$     $\mathcal{C}_{2}$$\mathcal{C}_{1}$$\mathcal{C}_{1}$$\mathcal{C}_{1}$$\mathcal{C}_{3}$
Figure 4: Proof of Proposition 5.
3.3 A reduction from SAT to $\psi$-NAN-dynamics
In this section, we explain how Theorem 1.c is obtained, and provide intermediate statements detailing a construction differing from the one presented in [10]. It will be beneficial in order to explain the variants later, and we provide full proofs.
To give an overview of this subsection, for a given MSO formula $\psi$ of rank $m$, whether $\Omega_{m}$ is sufficient (a model of $\psi$) or forbidden (a counter-model of $\psi$) indicates whether we respectively reduce from SAT ($\mathsf{NP}$}-hardness) or UNSAT ($\mathsf{coNP}$}-hardness). Given a propositional formula $S$ on $s$ variables, the configuration space $X$ is partitioned into $2^{s}$ subsets of configurations, so that $S$ is evaluated on each subset: